OpenStudy (anonymous):

Use L'Hospital's Rule to show that "Limit as X approaches infinity that (1+1/x)*x=e I dont know how to solve this...

OpenStudy (anonymous):

\[\lim_{X \rightarrow \infty}\left( 1+ 1/x \right)^x= e\] This is how the equation looks like. Its just I dont know how to take the derivative of this.

zepdrix (zepdrix):

Hmm so the first step is to make sure that it's of one of the specific indeterminate forms that we're allowed to apply L'Hospotal's to. Looks like when x heads towards infinity, our function is heading towards,\[\large 1^\infty\]While that is of indeterminate form, we need the form,\[\large\frac{0}{0} \quad \text{or}\quad \frac{\infty}{\infty}\] Another thing that might have helped to notice is that we can't take a derivative while there is a variable in the exponent AND base. So there are 2 reasons why we have to change the form around before we can differentiate.

OpenStudy (anonymous):

Okay yes I got \[1^\infty \] I just couldn't figure out how to take the derivative of the equation. Once I get that I know to just put infinity back in and see if it matches to one of the indeterminate forms.

zepdrix (zepdrix):

This next step can be a little confusing. Do you remember back to this identity?\[\large e^{\ln x}=x\]This is due to the fact that the exponential and the logarithm are inverse functions of one another. We're going to be applying this rule in reverse. We're STARTING with something involving x, and we're going to rewrite it as e^ln x.

zepdrix (zepdrix):

\[\huge \lim_{x \rightarrow \infty}\left(1+\frac{1}{x}\right)^x \quad \rightarrow \quad \lim_{x \rightarrow \infty}e^{\ln\left(\left(1+\frac{1}{x}\right)^x\right)}\]

zepdrix (zepdrix):

Understand how we did that? :O It's a little tricky.

OpenStudy (anonymous):

Yes, I do get that. :) Thank you! I forgot about the whole e^in of x=x so now that makes sense :) So now I just put infinity in for X and that should be my answer correct?

zepdrix (zepdrix):

I think we'll still have a problem, from here we wan to use rules of logarithms to move that X out of the log. Then we can get it into one of the 2 indeterminate forms that we want, and we'll be allowed to use L'Hop from there.

zepdrix (zepdrix):

From here, we just want to be working WITHIN the exponent. I know this is a little weird, but try to imagine yer floating in the sky, and the base e is below you. We'll just be dealing with this part of the problem,\[\huge \lim_{x \rightarrow \infty}{\ln\left(\left(1+\frac{1}{x}\right)^x\right)}\]

OpenStudy (anonymous):

Ohh okay! Lol I love the way you explain it btw :P

zepdrix (zepdrix):

Recall we can apply this rule,\[\large \log(b^a) \quad \rightarrow \quad a\cdot \log(b)\]Which will give us, Oh umm.. I'mma drop the limit for now, I don't want to keep writing it.\[\huge x\cdot \ln\left(1+\frac{1}{x}\right)\]

zepdrix (zepdrix):

Grrr gimme 2 minutes to finish my sammich >:O it's getting cold.. I knew I shouldn't had started a problem right now lol Sec brb :D

OpenStudy (anonymous):

Lol no problem :) Thanks :D

OpenStudy (anonymous):

Once we are done with this I actually have a couple other problems and if you wouldnt mind helping me with them, I would be very thankful :)

zepdrix (zepdrix):

If we were to let x approach infinity at this point, you would see that we have a problem approaching something like this,\[\large 0\cdot \ln(1)\quad \rightarrow \quad 0\cdot 0\]Hmm that's still no good, we need the QUOTIENT of functions to be able to apply L'Hop.

zepdrix (zepdrix):

Ahh that's frustrating, couldn't reconnect D:< Finally!

OpenStudy (anonymous):

So we will take the quotient of 1/X?

zepdrix (zepdrix):

This next step is a little tricky, we're gonna do something silly with the x term. We can rewrite it as the reciprocal of the reciprocal. We're sort of like.... flipping it, twice... which doesn't change the value of it. Lemme know if this is confusing or not.\[\huge x\cdot \ln\left(1+\frac{1}{x}\right) \quad \rightarrow \quad \frac{\ln\left(1+\frac{1}{x}\right)}{\frac{1}{x}}\]

zepdrix (zepdrix):

Ahh yer too fast >:O lol

OpenStudy (anonymous):

Okay....Ohh! I see! Then we would put infinity in and get something like \[\infty/infty\]

zepdrix (zepdrix):

Yah :) In this case, since x is approaching infinity, we get 0/0 which is also an acceptable form for L'Hop.

OpenStudy (anonymous):

Yay! Thank you!

zepdrix (zepdrix):

So apply L'Hop from this point, you'll get a nice cancellation. And then DON'T FORGET ABOUT THE E. Remember? We're floating in da sky! :O

OpenStudy (anonymous):

Okay what would I do with the E?

zepdrix (zepdrix):

Ignore it until the very very end. Do some L'Hop. Figure out what you're getting for a limit. Then as a final step, write it as e^ to that value.

OpenStudy (anonymous):

So I take the deriavtive of \[\ln \left( 1+1/x \right)/\left( 1/x \right)\]

zepdrix (zepdrix):

yah

zepdrix (zepdrix):

Well remember that you're taking the derivative of the top and bottom SEPARATELY. So you don't have an ugly quotient rule problem like it might look at first glance.

OpenStudy (anonymous):

Lol yeah thats what I thought at first but I remembered I was supposed to look at them SEPARATELY.

OpenStudy (anonymous):

Okay so how would I take the derivate of \[\ln \left( 1+1/x \right)\] I have the derrivate of 1/x which is just -1/x^2

zepdrix (zepdrix):

\[\huge \left(\ln x\right)'=\frac{1}{x}\]\[\huge \left(\ln\left(1+\frac{1}{x}\right)\right)'\quad =\quad \frac{1}{\left(1+\frac{1}{x}\right)}\left(1+\frac{1}{x}\right)'\]

zepdrix (zepdrix):

Don't forget the chain rule! :O That's what the prime term is on the right, have to differentiate that part still. Should give you something similar to the derivative of 1/x that you already took earlier ;U

OpenStudy (anonymous):

So now I put in infinity and I get 1? So its e^1?

zepdrix (zepdrix):

Yah looks good! :) Remember at the start what they were asking for? They actually gave you the answer, they said PROVE IT!! NOW!!!

OpenStudy (anonymous):

Lol wow they did :P Lol well thank you :D